OpenStudy (anonymous):

Find the dimensions of the rectangle of greatest area that can be inscribed in a semicircle of radius R.

OpenStudy (anonymous):

u look sad lol

OpenStudy (anonymous):

I can get part-way. Let the semi-circle be sitting on its diameter. Inscribe a rectangle of vertical height a, draw radius to the point where a strikes the circumference. use the triangle of the horizontal radius and a and the sloping radius, so that a = R sin @ half the base of the rectangle, along the diameter is w/2 (w/2) = R cos @ the area of the rectangle is a w = 2 R^2 sin @ cos @ and I would bet that this is maximized when @=45o, though I cannot prove it. area = a w = 2 R^2 (0.707)().707) = R^2. looks good, though speculative

OpenStudy (anonymous):

it says rectangle has to be inscribed in a semicircle !!

OpenStudy (anonymous):

Opps..my bad

OpenStudy (anonymous):

Yeah the rectangle is inside the circle lol

OpenStudy (anonymous):

Noooo rectangle must be INscribed, thus wholly within the semicircle. I do wish I could draw that well, though.

OpenStudy (anonymous):